### v2

# Compositional Risk Minimization

Abstract
Compositional generalization is a crucial step towards developing data-efficient intelligent machines that generalize in human-like ways. In this work, we tackle a challenging form of distribution shift, termed compositional shift, where some attribute combinations are completely absent at training but present in the test distribution. This shift tests the model’s ability to generalize compositionally to novel attribute combinations in discriminative tasks. We model the data with flexible additive energy distributions, where each energy term represents an attribute, and derive a simple alternative to empirical risk minimization termed compositional risk minimization (CRM). We first train an additive energy classifier to predict the multiple attributes and then adjust this classifier to tackle compositional shifts. We provide an extensive theoretical analysis of CRM, where we show that our proposal extrapolates to special affine hulls of seen attribute combinations. Empirical evaluations on benchmark datasets confirms the improved robustness of CRM compared to other methods from the literature designed to tackle various forms of subpopulation shifts.
[…]
Divyat Mahajan at
The ability to make sense of the rich complexity of the sensory world by decomposing it into sets of elementary factors and recomposing these factors in new ways is a hallmark of human intelligence. This capability is typically grouped under the umbrella term compositionality (Fodor and Pylyshyn, 1988; Montague, 1970). Compositionality underlies both semantic understanding and the imaginative prowess of humans, enabling robust generalization and extrapolation. For instance, human language allows us to imagine situations we have never seen before, such as “a blue elephant riding a bicycle on the Moon.” While most works on compositionality have focused on its generative aspect, i.e., imagination, as seen in diffusion models (Yang et al., 2023a), compositionality is equally important in discriminative tasks. In these tasks, the goal is to make predictions in novel circumstances that are best described as combinations of circumstances seen before. In this work, we dive into this less-explored realm of compositionality in discriminative tasks.
We work with multi-attribute data, where each input (e.g., an image) is associated with multiple categorical attributes, and the task is to predict an attribute or multiple attributes. During training, we observe inputs from only a subset of all possible combinations of individual attributes, and during test we will see novel combinations of attributes never seen at training. Following Liu et al. (2023), we refer to this distribution shift as compositional shift. Towards the goal of tackling these compositional shifts, we develop an adaptation of naive discriminative Empirical Risk Minimization (ERM) tailored for multi-attribute data under compositional shifts. We term our approach Compositional Risk Minimization (CRM). The foundations of CRM are built on additive energy distributions that are studied in generative compositionality (Liu et al., 2022a), where each energy term represents one attribute. In CRM, we first train an additive energy classifier to predict all the attributes jointly, and then we adjust this classifier for compositional shifts.
Figure 1: The additive energy classifier trained in CRM computes the logits for each group $z=(y,a)$ by adding the energy components of each attribute via boradcasting. For the train logits, we add the log of the prior probabilities and a learned bias $B(y,a)$ for the groups present in train data. At test time, the log prior term is replaced with the log of the test prior (if available, otherwise assumed to be uniform), and the biases for novel test groups, $B^{\star}(y,a)$, are extrapolated using Eq.11. Finally, we obtain $p(y,a|x)$ by applying softmax function on the adjusted logits. This adaptation from train to test is possible because of the additive energy distribution $p(x|y,a)$, which allows the model to factorize the distribution into distinct components associated with each attribute.
Our main contributions are as follows:
  • Theory of discriminative compositional shifts: For the family of additive energy distributions, we prove that additive energy classifiers generalize compositionally to novel combinations of attributes represented by a special mathematical object, which we call discrete affine hull. Our characterization of extrapolation is sharp, i.e., we show that it is not possible to generalize beyond discrete affine hull. We show that the volume of discrete affine hull grows very fast in the number of training attribute combinations thus generalizing to many attribute combinations. The proof techniques developed in this work are very different from existing works on distribution shifts and hence may be of independent interest.
  • A practical method: CRM is a simple algorithm for training classifiers, which first trains an additive energy classifier and then adjusts the trained classifier for tackling compositional shifts. We empirically validate the superiority of CRM to other methods previously proposed for addressing subpopulation shifts.
1 Problem Setting
1.1 Generalizing under Compositional Shifts
In compositional generalization, we aim to build a classifier that performs well in new contexts that are best described as a novel combination of seen contexts. Consider an input $x$ (e.g., image), this input belongs to a group that is characterized by an attribute vector $z=(z_{1},\ldots,z_{m})$ (e.g., class label, background label), where $z_{i}$ corresponds to the value of $i^{th}$ attribute. There are $m$ attributes and each attribute $z_{i}$ can take $d$ possible values. So $z\in\mathcal{Z}$ with $\mathcal{Z}=\{1,\ldots,d\}^{m}$.
[…]
We denote the support of each attribute component $z_{i}$ under training distribution as $\mathcal{Z}_{i}^{\mathsf{train}}$ and the support of $z$ under training distribution as $\mathcal{Z}^{\mathsf{train}}$. The corresponding supports for the test distribution are denoted as $\mathcal{Z}_{i}^{\mathsf{test}}$ and $\mathcal{Z}^{\mathsf{test}}$. We define the Cartesian product of marginal support under training as ${\mathcal{Z}^{\times}}\coloneq\mathcal{Z}_{1}^{\mathsf{train}}\times\mathcal{Z}_{2}^{\mathsf{train}}\times\cdots\mathcal{Z}_{m}^{\mathsf{train}}$.
In this work, we study compositional shifts from a training distribution $p$ to a test distribution $q$, characterized by:
  1. 1.
    $p(x|z)=q(x|z),\forall z\in{\mathcal{Z}^{\times}}.$
  2. 2.
    $\mathcal{Z}^{\mathsf{test}}\not\subseteq\mathcal{Z}^{\mathsf{train}}$ but $\mathcal{Z}^{\mathsf{test}}\subseteq{\mathcal{Z}^{\times}}$.
The first point states that the conditional density of inputs conditioned on attributes remains invariant from train to test, which can be understood as the data generation mechanism from attributes to the inputs remains invariant. What changes between train and test is thus due to only shifting prior probabilities of attributes from $p(z)$ to $q(z)$. The second point specifies how these differ in their support: at test we observe novel combinations of individual attributes but not a completely new individual attribute. The task of compositional generalization is then to build classifiers that are robust to such compositional distribution shifts. Also, we remark that the above notion should remind the reader of the notion of Cartesian Product Extrapolation (CPE) from Lachapelle et al. (2024). Specifically, if a model succeeds on test distributions $q(z)$ with support equal to the full Cartesian product ($\mathcal{Z}^{\mathsf{test}}={\mathcal{Z}^{\times}}$), then it is said to achieve CPE.
1.2 Additive Energy Distribution
We assume that $p(x|z)$ is of the form of an additive energy distribution (AED):
$p(x|z)=\frac{1}{\mathbb{Z}(z)}\exp\Big{(}-\sum_{i=1}^{m}E_{i}(x,z_{i})\Big{)}$ (1)
[…]
We thus have one energy term $E_{i}$ associated to each attribute $z_{i}$. Note that we do not make assumptions on $E_{i}$ except $\mathbb{Z}(z)<\infty$, leaving the resulting $p(x|z)$ very flexible. This form is a natural choice to model inputs that must satisfy a conjunction of characteristics (such as being a natural image of a landbird AND having a water background), corresponding to our attributes.
Recall $z=(z_{1},\ldots,z_{m})$ is a vector of $m$ categorical attributes that can each take $d$ possible values. We will denote as $\sigma(z)$ the representation of this attribute vector as a concatenation of $m$ one-hot vectors, i.e.
$\sigma(z)=[\mathrm{onehot}(z_{1}),\ldots,\mathrm{onehot}(z_{m})]^{\top}$
Thus $\sigma(z)$ will be a sparse vector of length $md$ containing $m$ ones. We also define a vector valued map $E(x)=[E_{1}(x,1),\ldots,E_{1}(x,d),\ldots,E_{m}(x,1),\ldots,E_{m}(x,d)]^{\top}$ where $E_{i}(x,z_{i})$ is the energy term for $i^{th}$ attribute taking the value $z_{i}$. This allows us to reexpress equation 1 using a simple dot product, denoted $\braket{\cdot,\cdot}$:
$p(x|z)=\frac{1}{\mathbb{Z}(z)}\exp\Big{(}-\braket{\sigma(z),E(x)}\Big{)},$ (2)
[…]
There are two lines of work that inspire the choice of additive energy distributions. Firstly, these distributions have been used to enhance compositionality in generative tasks  (Du et al., 2020, 2021; Liu et al., 2021) but they have not been used in discriminative compositionality. Secondly, for readers from the causal machine learning community, it may be useful to think of additive energy distributions from the perspective of the independent mechanisms principle (Janzing and Schölkopf, 2010; Parascandolo et al., 2018). The principle states that the data distribution is composed of independent data generation modules, where the notion of independence refers to algorithmic independence and not statistical independence. In these distributions, we think of energy function of an attribute as an independent function.
This is the right juncture to contrast AEDs with distributional assumptions in recent provable approaches to compositional generalization (Dong and Ma, 2022; Wiedemer et al., 2023, 2024; Brady et al., 2023; Lachapelle et al., 2024). These works assume labeling functions or decoders that are deterministic and additive over individual features, proving generalization over the Cartesian product of feature supports (further discussion in Appendix 7). While insightful, this assumption is restrictive, as each attribute combination corresponds to a single observation with limited generative interactions. In contrast, AEDs capture stochastic decoders, offering a more flexible way to model inputs as a conjunction of characteristics (see Appendix 8).
2 Provable Compositional Generalization
Our goal is to learn a model yielding a $\hat{q}(z|x)$ that will match the test distribution $q(z|x)$ and thus allow us to predict the attributes at test time in a Bayes optimal manner. If we successfully learn the distribution $q(z|x)$, then we can straightforwardly predict the individual attributes $q(z_{i}|x)$, e.g., the bird class in Waterbirds dataset, by marginalizing over the rest, e.g., the background in Waterbirds dataset. Observe that $q(z|x)$ differs from the training $p(z|x)$, which can be estimated through standard ERM with cross-entropy loss. Since some attributes $z$ observed at test time are never observed at train time, the distribution learned via ERM assigns a zero probability to these attributes and thus it cannot match the test distribution $q(z|x)$.
In what follows, we first introduce a novel mathematical object termed Discrete Affine Hull over the set of attributes. We then describe a generative approach for classification that requires us to learn $p(x|z)$ including the partition function, which is not practical. Next, we describe a purely discriminative approach that circumvents the issue of learning $\hat{p}(x|z)$ and achieves the same extrapolation guarantees as the generative approach. We present the generative approach as it allows to understand the results more easily. Building generative models based on our theory is out of scope of this work but is an exciting future work.
2.1 Discrete Affine Hull
We define the discrete affine hull of a set of attribute vectors $\mathcal{A}=\{z^{(1)},\ldots,z^{(k)}\}$ where $z^{(i)}\in\mathcal{Z}$ as:
$\mathsf{DAff}(\mathcal{A})=\Big{\{}z\in\mathcal{Z}\;|\;\exists\;\alpha\in\mathbb{R}^{k},\sigma(z)=\sum_{i=1}^{k}\alpha_{i}\sigma\big{(}z^{(i)}\big{)},\sum_{i=1}^{k}\alpha_{i}=1\Big{\}}$
In other words, the discrete affine hull of $\mathcal{A}$ consists of all attribute vectors whose one-hot encoding lies in the (regular) affine hull of the one-hot encodings of the attribute vectors of $\mathcal{A}$. This construct helps characterize which new attribute combinations we can extrapolate to.
As an illustration, consider the Waterbirds dataset, where we observe three of four possible groups. In one-hot encoding, WB is $[1,0]$, LB is $[0,1]$, Water is $[1,0]$, and Land is $[0,1]$. We show that the missing attribute vector WB on L, represented as [1 0 0 1]delimited-[]1001[1\;0\;0\;1][ 1 0 0 1 ], can be expressed as an affine combination of the observed vectors, meaning the discrete affine hull of three one-hot concatenated vectors contains all four possible combinations.
$(+1)\cdot\begin{bmatrix}0\\ 1\\ 0\\ 1\end{bmatrix}\;\;+\;\;(-1)\cdot\begin{bmatrix}0\\ 1\\ 1\\ 0\end{bmatrix}\;\;+\;\;(+1)\cdot\begin{bmatrix}1\\ 0\\ 1\\ 0\end{bmatrix}=\begin{bmatrix}1\\ 0\\ 0\\ 1\end{bmatrix}$ (3)
In Section 10.5, we generalize this finding to a formal mathematical characterization of discrete affine hulls, providing a visualization method. In Section 10.7, we show how discrete affine hulls generalize the extrapolation of additive functions studied in Dong and Ma (2022); Lachapelle et al. (2024) over discrete domains. We also show how our results lead to a sharp characterization of extrapolation of these functions. Throughout, “affine hull” refers to the discrete affine hull.
2.2 Extrapolation of Conditional Density
[…]
Theorem 1.
If the true and learned distribution ($p(\cdot|z)$ and $\hat{p}(\cdot|z)$) are AED, then $\hat{p}(\cdot|z)=p(\cdot|z),\forall z\in\mathcal{Z}^{\mathsf{train}}\implies\hat{p}(\cdot|z^{\prime})=p(\cdot|z^{\prime}),\forall z^{\prime}\in\mathsf{DAff}(\mathcal{Z}^{\mathsf{train}})$.
The result above argues that so long as the group $z^{\prime}$ is in the discrete affine hull of $\mathcal{Z}^{\mathsf{train}}$, the estimated density extrapolates to it. We provide a proof sketch ahead, with the complete proof in Section 10.1.
Proof sketch: Under perfect maximum likelihood maximization $\hat{p}(x|z)=p(x|z),\forall z\in\mathcal{Z}^{\mathsf{train}}$. Replacing these densities by their expressions and taking their $\log$ we obtain
$\braket{\sigma(z),\hat{E}(x)}=\braket{\sigma(z),E(x)}+C(z),\forall z\in\mathcal{Z}^{\mathsf{train}}$ (4)
where $C(z)=\log\big{(}\mathbb{Z}(z)/\hat{\mathbb{Z}}(z)\big{)}$.
For any $z^{\prime}\in\mathsf{DAff}(\mathcal{Z}^{\mathsf{train}})$, by definition there exists $\alpha$ such that $\sigma(z^{\prime})=\sum_{z\in\mathcal{Z}^{\mathsf{train}}}\alpha_{z}\sigma(z)$. Thus $\braket{\sigma(z^{\prime}),\hat{E}(x)}=\sum_{z\in\mathcal{Z}^{\mathsf{train}}}\alpha_{z}\braket{\sigma(z),\hat{E}(x)}$, by linearity of the dot product. Substituting the expression for $\braket{\sigma(z),\hat{E}(x)}$ from equation 4, this becomes
$\begin{split}\braket{\sigma(z^{\prime}),\hat{E}(x)}=\sum_{z\in\mathcal{Z}^{\mathsf{train}}}\alpha_{z}\Big{(}\braket{\sigma(z),E(x)}+C(z)\Big{)}\\ =\braket{\sigma(z^{\prime}),E(x)}+\sum_{z\in\mathcal{Z}^{\mathsf{train}}}\alpha_{z}C(z),\end{split}$ (5)
From equation 5, we can conclude that $\braket{\sigma(z^{\prime}),\hat{E}(x)}$ estimates $\braket{\sigma(z^{\prime}),E(x)}$ perfectly up to a constant error that does not depend on $x$. This difference of constant is absorbed by the partition function and hence the conditional densities match: $\hat{p}(x|z^{\prime})=p(x|z^{\prime})$.
Classifier based on conditional density $p(x|z)$.
If, on data from training distribution $p$, we were able to train a good conditional density estimate $\hat{p}(x|z),\forall z\in\mathcal{Z}^{\mathsf{train}}$, then Theorem 1 implies that $\hat{p}(x|z^{\prime})$ will also be a good estimate of $p(x|z^{\prime})$ for new unseen attributes $z^{\prime}\in\mathsf{DAff}(\mathcal{Z}^{\mathsf{train}})$. Provided $\mathcal{Z}^{\mathsf{test}}\subseteq\mathsf{DAff}(\mathcal{Z}^{\mathsf{train}})$, it is then straightforward to obtain a classifier that generalizes to compositionally-shifted test distribution $q$. Indeed, we have
$q(z^{\prime}|x)$ $=\frac{q(x|z^{\prime})q(z^{\prime})}{\sum_{z^{\prime\prime}\in\mathcal{Z}^{\mathsf{test}}}q(x|z^{\prime\prime})q(z^{\prime\prime})}$ (6)
$=\frac{p(x|z^{\prime})q(z^{\prime})}{\sum_{z^{\prime\prime}\in\mathcal{Z}^{\mathsf{test}}}p(x|z^{\prime\prime})q(z^{\prime\prime})}\approx\frac{\hat{p}(x|z^{\prime})q(z^{\prime})}{\sum_{z^{\prime\prime}\in\mathcal{Z}^{\mathsf{test}}}\hat{p}(x|z^{\prime\prime})q(z^{\prime\prime})}$
where we used the property of compositional shifts $q(x|z)=p(x|z)$. If we know test group prior $q(z^{\prime})$ (or e.g. assume it to be uniform), we can directly use the expression in RHS to correctly compute the test group probabilities $q(z|x)$, even for groups never seen at training.
2.3 Extrapolation of Discriminative Model
[…]
But is it really necessary to precisely model the conditional density of high dimensional $x$, when our goal is simply to predict a few classes and attributes $z$, given $x$? We will now develop an alternative approach, Compositional Risk Minimization (CRM), that achieves a similar extrapolation result as Theorem 1, while being based on simple discriminative classifier training. It sidesteps the need and difficulties of explicitly modeling $p(x|z)$ and doesn’t require dealing with the partition function throughout training.
Observe that if we apply Bayes rule to the AED $p(x|z)$ in equation 2, we get
$p(z|x)=\frac{p(x|z)p(z)}{\sum_{z^{\prime}\in\mathcal{Z}^{\mathsf{train}}}p(x|z^{\prime})p(z^{\prime})}$
$=\frac{\exp\Big{(}-\braket{\sigma(z),E(x)}+\log p(z)-\log\mathbb{Z}(z)\Big{)}}{\sum_{z^{\prime}\in\mathcal{Z}^{\mathsf{train}}}\exp\Big{(}-\braket{\sigma(z^{\prime}),E(x)}+\log p(z^{\prime})-\log\mathbb{Z}(z^{\prime})\Big{)}}$
We thus define our additive energy classifier as follows. To guarantee that we can model this $p(z|x)$, we use a model with the same form. For each $z\in\mathcal{Z}^{\mathsf{train}}$
$\tilde{p}(z|x)=\frac{\exp\Big{(}-\braket{\sigma(z),\tilde{E}(x)}+\log\hat{p}(z)-\tilde{B}(z)\Big{)}}{\sum_{z^{\prime}\in\mathcal{Z}^{\mathsf{train}}}\exp\Big{(}-\braket{\sigma(z^{\prime}),\tilde{E}(x)}+\log\hat{p}(z^{\prime})-\tilde{B}(z^{\prime})\Big{)}}$ (7)
where $\hat{p}(z)$ is the empirical estimate of the prior over $z$, i.e., $p(z)$, $\tilde{E}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{md}$ is a function to be learned, bias $\tilde{B}$ is a lookup table containing a learnable offset for each combination of attribute. Given a data point $(x,z)$, loss $\ell(z,\tilde{p}(\cdot|x))=-\log\tilde{p}(z|x)$ measures the prediction performance of $\tilde{p}(\cdot|x)$. The risk, defined as the expected loss, corresponds to the negated conditional log-likelihood:
$R(\tilde{p})=\mathbb{E}_{(x,z)\sim p}\big{[}\ell(z,\tilde{p}(\cdot|x))\big{]}=\mathbb{E}_{(x,z)\sim p}\big{[}-\log\tilde{p}(z|x)\big{]}$ (8)
In the first step of CRM, we minimize the risk $R$.
$\hat{E},\hat{B}\in\operatorname*{arg\,min}_{\tilde{E},\tilde{B}}R(\tilde{p})$ (9)
If the minimization is over arbitrary functions, then $\hat{p}(\cdot|x)=p(\cdot|x),\forall x\in\mathbb{R}^{n}$. In the second step of CRM, we compute our final predictor $\hat{q}(z|x)$ as follows. Let $\hat{q}(z)$ be an estimate of the marginal distribution over the attributes $q(z)$ with support $\hat{\mathcal{Z}}^{\mathsf{test}}$. For each $z\in\mathcal{Z}^{\mathsf{test}}$
$\hat{q}(z|x)=\frac{\exp\Big{(}-\braket{\sigma(z),\hat{E}(x)}+\log\hat{q}(z)-B^{\star}(z)\Big{)}}{\sum_{z^{\prime}\in\hat{\mathcal{Z}}^{\mathsf{test}}}\exp\Big{(}-\braket{\sigma(z^{\prime}),\hat{E}(x)}+\log\hat{q}(z^{\prime})-B^{\star}(z^{\prime})\Big{)}}$ (10)
where, $B^{\star}$ is the extrapolated bias defined as
$B^{\star}(z)=\log\mathbb{E}_{x\sim p}\Big{[}\frac{\exp\Big{(}-\braket{\sigma(z),\hat{E}(x)}\Big{)}}{\sum_{\tilde{z}\in\mathcal{Z}^{\mathsf{train}}}\exp\Big{(}-\braket{\sigma(\tilde{z}),\hat{E}(x)}+\log p(\tilde{z})-\hat{B}(\tilde{z})\Big{)}}\Big{]}$ (11)
where $\hat{E}$, $\hat{B}$ are the solutions from optimization equation 9. Note that $\hat{B}(z)$ was learned for all $z\in\mathcal{Z}^{\mathrm{train}}$ but never for $z\in\mathcal{Z}^{\mathrm{test}}$, hence the necessity of extrapolation $B^{*}$. Each of these steps is easy to operationalize. We explain the process and provide pseudocode in Section 3.
Theorem 2.
Consider the setting where $p(.|z)$ follows AED $\forall z\in{\mathcal{Z}^{\times}}$, the test distribution $q$ satisfies compositional shift characterization and $\mathcal{Z}^{\mathsf{test}}\subseteq\mathsf{DAff}(\mathcal{Z}^{\mathsf{train}})$. If $\hat{p}(z|x)=p(z|x),\forall z\in\mathcal{Z}^{\mathsf{train}},\forall x\in\mathbb{R}^{n}$ and $\hat{q}(z)=q(z),\forall z\in\mathcal{Z}^{\mathsf{test}}$, then the output of CRM (equation 10) matches the test distribution, i.e., $\hat{q}(z|x)=q(z|x),\forall z\in\mathcal{Z}^{\mathsf{test}},\forall x\in\mathbb{R}^{n}$.
A complete proof is provided in Section 10.2. Observe that $\hat{p}(\cdot|x)=p(\cdot|x)$ is a condition that even a model trained via ERM can satisfy (with sufficient capacity and data) but it cannot match the true $q(\cdot|x)$. In contrast, CRM optimally adjusts the additive-energy classifier for the compositional shifts. CRM requires the knowledge of test prior $q(z)$ but the choice of uniform distribution over all possible groups is a reasonable one to make in the absence of further knowledge. Notice how learned bias $\hat{B}(z)$ can only be fitted for $z\in\mathcal{Z}^{\mathrm{train}}$, remaining undefined for $z^{\prime}\notin\mathcal{Z}^{\mathrm{train}}$. But we can compute the extrapolated bias $B^{\star}(z^{\prime})$, $\forall z^{\prime}\in\mathcal{Z}^{\mathrm{test}}$, based remarkably on only data from the train distribution.
Illustrating CRM’s adaptation to test distribution.
To better convey how CRM can adapt to the Bayes optimal classifier of the test distribution, we provide an example. Consider a two-dimensional setting, where the distribution of $x\in\mathbb{R}^{2}$ conditioned on the attributes $z_{1}\in\{-1,1\}$ and $z_{2}\in\{-1,1\}$ is a Gaussian with mean $(z_{1},z_{2})$ and identity covariance. Suppose the training groups are drawn with equal probability and can take one of the following three possible values $(+1,+1),(-1,+1),(+1,-1)$. We do not observe data from the group $(-1,-1)$ during training, but at test time we draw samples from all the four groups with equal probability. First, we can show that the above distribution can be expressed as an additive energy distribution, as $E(x,(z_{1},z_{2}))=\frac{1}{4}\|x-(2z_{1},0)\|^{2}+\frac{1}{4}\|x-(0,2z_{2})\|^{2})$. For the task of classifying groups, the Bayes optimal classifier has a closed form solution, where each decision region is an intersection of two half-spaces. Figure 2 shows that CRM learns the Bayes optimal classifier on the training distribution, enables shifting from train prior to test prior to yield the Bayes optimal classifier for the test distribution, and correctly generalize to the unseen $(-1-1)$ group. For further details, including illustration of the failure of ERM-trained binary classifier on this problem, see Section 11. Beyond this toy example, we highlight that the additive energy form supports modeling nearly arbitrarily complex distributions.
Figure 2: Extrapolating to an unseen test group. The distribution to model corresponds to a mixture of 4 Gaussians. But the $(-1,-1)$ group (pink dashed) has zero prior probability in the training distribution, i.e., is absent from the training set. As a result, discriminative training (left) learns only three decision regions and would misclassify test points from $(-1,-1)$. CRM adjusts the prior at test time (right) to a uniform distribution over all attribute combinations, enabling it to recover four decision regions and correctly generalize to the unseen $(-1,-1)$ group. The additive energy form composes the attributes’ information to yield this unseen group’s location. Decision regions were obtained from finite-data simulations, leading to minor imperfections.
Analyzing growth of Discrete Affine Hull.
In the discussion so far, we have relied on a crucial assumption that the attribute combinations in the test distribution are in the affine hull. Is this also a necessary condition? Can we generalize to attributes outside the affine hull? We consider the task of learning $p(\cdot|z)$ from Theorem 1 and the task of learning $q(\cdot|x)$ from Theorem 2. In  Section 10.6, we show that the restriction to affine hulls is indeed necessary. Under the assumption of compositional shifts $\mathcal{Z}^{\mathsf{test}}$ is only restricted to be a subset of the Cartesian product set ${\mathcal{Z}^{\times}}$, but our results so far have required us to restrict the support further by confining it to the affine hull, i.e., $\mathcal{Z}^{\mathsf{test}}\subseteq\mathsf{DAff}(\mathcal{Z}^{\mathsf{train}})\subseteq{\mathcal{Z}^{\times}}$. This leads us to a natural question. If the training groups that form $\mathcal{Z}^{\mathsf{train}}$ are drawn at random, then how many groups do we need such that the affine hull captures ${\mathcal{Z}^{\times}}$, i.e., $\mathsf{DAff}(\mathcal{Z}^{\mathsf{train}})={\mathcal{Z}^{\times}}$, at which point CRM can achieve Cartesian Product Extrapolation (CPE). Another way to think about this is to say, how fast does the affine hull grow and capture the Cartesian product set ${\mathcal{Z}^{\times}}$? Next, we answer this question.
Consider the the general setting with $m$ attributes, where each attribute takes $d$ possible values, leading to $d^{m}$ possible attribute combinations. Suppose we sample $s$ attribute vectors $z$ that comprise the support $\mathcal{Z}^{\mathsf{train}}$ uniformly at random (with replacement) from these $d^{m}$ possibilites. In the next theorem, we show that if the number of sampled attribute vectors exceeds $2c(md+d\log(d))$, then $\mathsf{DAff}(\mathcal{Z}^{\mathsf{train}})$ contains all the possible $d^{m}$ combinations with a high probability (greater than $1-\frac{1}{c}$), hence CRM achieves CPE. We want to emphasize this surprising finding: with almost a linear growth in $m$ and $d$, CRM generalizes to exponentially many $d^{m}$ groups.
Theorem 3.
Consider the setting where $p(.|z)$ follows AED $\forall z\in{\mathcal{Z}^{\times}}$, $\mathcal{Z}^{\mathsf{train}}$ comprises of $s$ attribute vectors $z$ drawn uniformly at random from ${\mathcal{Z}^{\times}}$, and the test distribution $q$ satisfies compositional shift characterization with $\mathcal{Z}^{\mathsf{test}}=\mathcal{Z}^{\times}$. If $s\geq 2c(md+d\log(d))$, where $d$ is sufficiently large, $\hat{p}(z|x)=p(z|x),\forall z\in\mathcal{Z}^{\mathsf{train}},\forall x\in\mathbb{R}^{n}$, $\hat{q}(z)=q(z),\forall z\in{\mathcal{Z}^{\times}}$, then the output of CRM (equation 10) matches the test distribution, i.e., $\hat{q}(z|x)=q(z|x)$, $\forall z\in{\mathcal{Z}^{\times}},\forall x\in\mathbb{R}^{n}$, with probability greater than $1-\frac{1}{c}$.
In Section 10.3, we first present the proof for the case with $m=2$ attributes and provide visual illustrations to assist the reader. It is then followed by the more involved proof for the general case of $m$ attributes in Section 10.4.
3 Algorithm for CRM
Input: Training set $\mathcal{D}^{\mathrm{train}}=\{(x,z)\}$
Output: Classifier parameters $\hat{\theta}$, $\hat{W}$, $B^{\star}$
Training:
  • Estimate train prior $\hat{p}(z)$ based on group counts in $\mathcal{D}^{\mathrm{train}}$
  • Compute the energy terms as $\tilde{E}(x)=\tilde{W}\phi(x;\tilde{\theta})$
  • CRM Step 1: Train additive energy classifier $\tilde{p}$ (e.q. 7) by emprirical risk minimization:
    $\hat{\theta},\hat{W},\hat{B}\in\operatorname*{arg\,min}_{\tilde{\theta},\tilde{W},\tilde{B}}R(\tilde{p})$
  • CRM Step 2: Estimate extrapolated bias $B^{\star}$ (e.q. 11) via an average over training examples.
Inference on test point $x$:
  • Set $\hat{q}(z)$ as uniform prior over all groups.
  • Compute test group probabilities $\hat{q}(z|x)$ via e.q. 10 using $\hat{E}$ and $B^{\star}$ learned during training.
Algorithm 1 Compositional Risk Minimization (CRM)
In a nutshell, CRM consists of: a) training additive energy classifier $\hat{p}(z|x)$ (e.q. 7) by maximum likelihood (e.q.9) for trainset group prediction; b) compute extrapolated biases $B^{\star}$ (e.q.11); c) infer group probabilities on compositionally shifted test distribution using $\hat{q}(z|x)$ (e.q 10). Algorithm 1 provides the associated pseudo-code, where we have a basic architecture using a deep network backbone $\phi(x;\theta)$ followed by a linear mapping (matrix $W$)  [^1] Instead of linear, we could use separate non-linear heads to obtain the energy components for each attribute.. For the case where we have 2 attributes $z=(y,a)$ (illustrated in Figure 1), we provide a detailed algorithm and PyTorch implementation in Section 9.
4 Experiments
4.1 Setup
We evaluate CRM on widely recognized benchmarks for subpopulation shifts (Yang et al., 2023b), that have attributes $z=(y,a)$, where $y$ denotes the class label and $a$ denotes the spurious attribute ($y$ and $a$ are correlated). However, the standard split between train and test data mandated in these benchmarks does not actually evaluate compositional generalization capabilities, because both train and test datasets contain all the groups ($\mathcal{Z}^{\mathsf{train}}=\mathcal{Z}^{\mathsf{test}}={\mathcal{Z}^{\times}}$). Therefore, we repurpose these benchmarks for compositional shifts by discarding samples from one of the groups ($z$) in the train (and validation) dataset; but we don’t change the test dataset, i.e., $z\not\in\mathcal{Z}^{\mathsf{train}}$ but $z\in\mathcal{Z}^{\mathsf{test}}$. Let us denote the data splits from the standard benchmarks as $({\mathcal{D}}_{\mathsf{train}},{\mathcal{D}}_{\mathsf{val}},{\mathcal{D}}_{\mathsf{test}})$. Then we generate multiple variants of compositional shifts $\{({\mathcal{D}}_{\mathsf{train}}^{\neg z},{\mathcal{D}}_{\mathsf{val}}^{\neg z},{\mathcal{D}}_{\mathsf{test}})\;|\;z\in{\mathcal{Z}^{\times}}\}$, where ${\mathcal{D}}_{\mathsf{train}}^{\neg z}$ and ${\mathcal{D}}_{\mathsf{val}}^{\neg z}$ are generated by discarding samples from ${\mathcal{D}}_{\mathsf{train}}$ and ${\mathcal{D}}_{\mathsf{val}}$ that belong to the group $z$.
Following this procedure, we adapted Waterbirds (Wah et al., 2011), CelebA (Liu et al., 2015), MetaShift (Liang and Zou, 2022), MultiNLI (Williams et al., 2017), and CivilComments Borkan et al. (2019) for experiments. We also experiment with the NICO++ dataset (Zhang et al., 2023), where we already have $\mathcal{Z}^{\mathsf{train}}\subsetneq\mathcal{Z}^{\mathsf{test}}={\mathcal{Z}^{\times}}$ as some groups were not present in the train dataset. However, these groups are still present in the validation dataset ($\mathcal{Z}^{\mathsf{val}}={\mathcal{Z}^{\times}}$). Hence, the only transformation we apply to NICO++ is to drop samples from the validation dataset so that $\mathcal{Z}^{\mathsf{train}}=\mathcal{Z}^{\mathsf{val}}$. Note that our benchmarks cover diverse scenarios, with binary (Waterbirds, CelebA, MetaShift) and non-binary attributes (MultiNLI, CivilComments, NICO++), resulting in total groups varying from $4$ to $360$ (NICO++).
For baselines, we train classifiers via ERM, Group Distributionally Robust Optimization (GroupDRO) (Sagawa et al., 2019), Logit Correction (LC) (Liu et al., 2022b), supervised logit adjustment (sLA) (Tsirigotis et al., 2024), Invariant Risk Minimization (IRM) (Arjovsky et al., 2019), Risk Extrapolation (VREx) (Krueger et al., 2021), and Mixup (Zhang et al., 2017). In all cases we employ a pretrained architecture as the representation network $\phi$, followed by a linear layer $W$ to get class predictions, and fine-tune them jointly (see Section 12.3 for details).
For evaluation metrics, we computed the average accuracy, group-balanced accuracy, and worst-group accuracy (WGA) on the test set. Due to imbalances in group distribution, a method can obtain good average accuracy despite having bad worst-group accuracy. Therefore, WGA is more indicative of robustness to spurious correlations (Section 12.2).
4.2 Results
Table 1 shows the results of our experiment, where due to space constraints, we only compare with the best performing baselines and don’t report the group balanced accuracy. Complete results with comparisons with IRM, VREx, and Mixup are provided in Appendix 13.1 (Table 5). For each dataset and metric, we report the average performance over its various compositional shift scenarios $\{({\mathcal{D}}_{\mathsf{train}}^{\neg z},{\mathcal{D}}_{\mathsf{val}}^{\neg z},{\mathcal{D}}_{\mathsf{test}})\;|\;z\in{\mathcal{Z}^{\times}}\}$ (detailed results for all scenarios are in Section 13.2). In all cases, CRM either outperforms or is competitive with the baselines in terms of worst group accuracy (WGA).
Further, for Waterbirds and MultiNLI, while the logit adjustment baselines appear competitive with CRM on average, if we look more closely at the worst case compositional shift scenario, we find these baselines fare much worse than CRM. For Waterbirds, LC obtains $69.0\%$ WGA while CRM obtains $73.0\%$ WGA for the worst case scenario of dropping the group $(0,1)$ (Table 6). Similarly, for MultiNLI, sLA obtains $19.7\%$ WGA while CRM obtains $31.0\%$ WGA for the worst case scenario of dropping the group $(0,0)$ (Table 9).
We also report the worst group accuracy (other metrics in Table 14Section 13.5) for the original benchmark $({\mathcal{D}}_{\mathsf{train}},{\mathcal{D}}_{\mathsf{val}},{\mathcal{D}}_{\mathsf{train}})$, which was not transformed for compositional shifts, denoted WGA (No Groups Dropped). This can be interpreted as the “oracle” performance for that benchmark, and we can compare methods based on the performance drop in WGA due to discarding groups in compositional shifts. ERM and GroupDRO appear the most sensitive to compositional shifts, and the logit adjustment baselines also show a sharp drop for the CelebA benchmark; while CRM is more robust to compositional shifts.
Multiple spurious attributes case. We benchmark CRM for class label prediction with multiple spurious attributes. For this, we augment the CelebA benchmark to have three additional attributes (total $m=5$ attributes). We provide setup details in Section 13.3, and Table 12 presents the associated results. We find that CRM still continues to be the superior method (as per WGA), in fact, the difference in WGA between CRM and runner up baseline increases to $28.9\%$, as compared to $14\%$ for the case of CelebA 2-attribute (Table 2). Hence, the benefits with CRM become more pronounced for multiple spurious attributes settings.
Importance of extrapolating the bias. We conduct an ablation study for CRM where we test a variant that uses the learned bias $\hat{B}$ (e.q. 9) instead of the extrapolated bias $B^{\star}$ (e.q. 11). Results are presented in Table 2. They show a significant drop in worst-group accuracy if we use the learned bias instead of the extrapolated one. Hence, our theoretically grounded bias extrapolation step is crucial to generalization under compositional shifts. In Section 13.4 (Table 13) we conduct further ablation studies, showing the impact of different choices of the test log prior.
Further, we analyze CRM’s performance with varying number of train groups, details in Section 13.6.
Dataset Method Average Acc WGA WGA (No Groups Dropped)
Waterbirds ERM $77.9$ $(0.1)$ $43.0$ $(0.2)$ $62.3$ $(1.2)$
G-DRO $77.9$ $(0.9)$ $42.3$ $(2.6)$ $87.3$ $(0.3)$
LC $88.3$ $(0.9)$ $75.5$ $(1.8)$ $88.7$ $(0.3)$
sLA $89.3$ $(0.4)$ $77.3$ $(1.4)$ $89.7$ $(0.3)$
CRM $87.1$ $(0.7)$ $78.7$ $(1.0)$ $86.0$ $(0.6)$
CelebA ERM $85.8$ $(0.3)$ $39.0$ $(0.3)$ $52.0$ $(1.0)$
G-DRO $89.2$ $(0.5)$ $67.8$ $(0.8)$ $91.0$ $(0.6)$
LC $91.1$ $(0.2)$ $57.4$ $(0.5)$ $90.0$ $(0.6)$
sLA $90.9$ $(0.2)$ $57.4$ $(1.3)$ $86.7$ $(1.9)$
CRM $91.1$ $(0.2)$ $81.8$ $(0.5)$ $89.0$ $(0.6)$
MetaShift ERM $85.7$ $(0.4)$ $60.5$ $(0.5)$ $63.0$ $(0.0)$
G-DRO $86.0$ $(0.3)$ $63.8$ $(1.1)$ $80.7$ $(1.3)$
LC $88.5$ $(0.0)$ $68.2$ $(0.5)$ $80.0$ $(1.2)$
sLA $88.4$ $(0.1)$ $63.0$ $(0.5)$ $80.0$ $(1.2)$
CRM $87.6$ $(0.3)$ $73.4$ $(0.4)$ $74.7$ $(1.5)$
MultiNLI ERM $68.4$ $(2.1)$ $7.5$ $(1.3)$ $68.0$ $(1.7)$
G-DRO $70.4$ $(0.2)$ $34.3$ $(0.2)$ $57.0$ $(2.3)$
LC $75.9$ $(0.1)$ $54.3$ $(1.0)$ $74.3$ $(1.2)$
sLA $76.4$ $(0.3)$ $55.0$ $(1.5)$ $71.7$ $(0.3)$
CRM $74.3$ $(0.3)$ $58.7$ $(1.4)$ $74.7$ $(1.3)$
CivilComments ERM $80.4$ $(0.2)$ $55.9$ $(0.2)$ $61.0$ $(2.5)$
G-DRO $80.1$ $(0.1)$ $61.6$ $(0.5)$ $64.7$ $(1.5)$
LC $80.7$ $(0.1)$ $65.7$ $(0.5)$ $67.3$ $(0.3)$
sLA $80.6$ $(0.1)$ $65.6$ $(0.2)$ $66.3$ $(0.9)$
CRM $83.7$ $(0.1)$ $67.9$ $(0.5)$ $70.0$ $(0.6)$
NICO++ ERM $85.0$ $(0.0)$ $35.3$ $(2.3)$ $35.3$ $(2.3)$
G-DRO $84.0$ $(0.0)$ $36.7$ $(0.7)$ $33.7$ $(1.2)$
LC $85.0$ $(0.0)$ $35.3$ $(2.3)$ $35.3$ $(2.3)$
sLA $85.0$ $(0.0)$ $33.0$ $(0.0)$ $35.3$ $(2.3)$
CRM $84.7$ $(0.3)$ $40.3$ $(4.3)$ $39.0$ $(3.2)$
Table 1: Robustness under compositional shift. We compare the proposed CRM method to baseline ERM classifier training with no group information, and to robust methods that leverage group labels: G-DRO, LC, and sLA. We report test Average Accuracy and Worst Group Accuracy (WGA) (mean over 3 random seeds), averaged as a group is dropped from training and validation sets. Last column is WGA under the dataset’s standard subpopulation shift benchmark, i.e. with no group dropped. All methods have a harder time to generalize when groups are absent from training, but CRM appears consistently more robust. Full results with balanced group accuracy, and baselines IRM, VREx, and Mixup in Appendix (Table 5).
Method Waterbirds CelebA MetaShift MulitNLI CivilComments NICO++
CRM ($\hat{B}$) $55.7$ $(1.0)$ $58.9$ $(0.4)$ $58.7$ $(0.6)$ $29.2$ $(2.1)$ $51.9$ $(1.0)$ $31.0$ $(1.0)$
CRM $78.7$ $(1.6)$ $81.8$ $(1.2)$ $73.4$ 0.7)0.7)0.7 ) $57.7$ $(3.0)$ $68.1$ $(0.5)$ $40.3$ $(4.3)$
Table 2: Importance of bias extrapolation. We report Worst Group Accuracy, averaged as a group is dropped from training and validation (standard error based on 3 random seeds). CRM $(\hat{B})$ is an ablated version of CRM where we use the trained bias $\hat{B}$ instead of the extrapolated bias $B^{\star}$ mandated by our theory. The extrapolation step appears crucial for robust compositional generalization. Merely adjusting logits based on shifting group prior probabilities does not suffice.
5 Related works
We briefly describe the relevant prior works, with a detailed discussion in Section 7.
Compositional Generalization Compositionality has long been seen as an essential capability (Fodor and Pylyshyn, 1988; Hinton, 1990; Plate et al., 1991; Montague, 1970) on the path to building human-level intelligence. The history of compositionality being too long to cover in detail here, we refer the reader to these surveys (Lin et al., 2023; Sinha et al., 2024). Most prior works have focused on generative aspect of compositionality, where the model needs to recombine individual distinct factors/concepts and generate the final output in the form of text (Gordon et al., 2019; Lake and Baroni, 2023) or image (Liu et al., 2022a; Wang et al., 2024). For image generation in particular, a fruitful line of work is rooted in additive energy based models (Du et al., 2020, 2021; Liu et al., 2021; Nie et al., 2021), which translates naturally to additive diffusion models (Liu et al., 2022a; Su et al., 2024). Our present work also leverages an additive energy form, but our focus is on learning classifiers robust under compositional shifts, rather than generative models.
Domain Generalization Generalization under subpopulation shifts, where certain groups or combinations of attributes are underrepresented in the training data, is a well-known challenge in machine learning. GroupDRO (Sagawa et al., 2019) is a prominent method that minimizes the worst-case group loss to improve robustness across groups. IRM (Arjovsky et al., 2019) encourages the model to learn invariant representations that perform well across multiple environments. Closely related to our proposed method are the logit adjustment methods, LC (Liu et al., 2022b) and sLA (Tsirigotis et al., 2024) that use logit adjustment for group robustness. There are many other interesting approaches that were proposed, see the survey (Zhou et al., 2022) for details. The theoretical guarantees developed for these approaches (Arjovsky et al., 2019; Rosenfeld et al., 2020; Ahuja et al., 2020) require a large diversity in terms of the environments seen at the training time. In our setting, we incorporate inductive biases based on additive energy distributions that help us arrive at provable generalization with limited diversity in the environments.
6 Conclusion
We provide a novel approach (CRM) based on flexible additive energy models for compositionality in discriminative tasks. CRM can provably extrapolate to novel attribute combinations within the discrete affine hull of the training support, where the affine hull grows quickly with the training groups to cover the Cartesian product extension of the training support. Our empirical results demonstrate that the additive energy assumption is sufficiently flexible to yield good classifiers for high-dimensional images, and CRM is able to extrapolate to novel combinations in $\mathsf{DAff}(\mathcal{Z}^{\mathsf{train}})$, without having to model high-dimensional $p(x|z)$ nor having to estimate their partition function. CRM is a simple and efficient algorithm that empirically proved consistently more robust to compositional shifts than approaches based on other logit-shifting schemes and GroupDRO.
Acknowledgements
We thank Yash Sharma for his contribution to early exploration of compositional generalization from a generative perspective, and Vinayak Tantia for having many years ago helped shed light on the challenges posed by compositionality in discriminative training. This research was entirely funded by Meta, in the context of Meta’s AI Mentorship program with Mila. Ioannis Mitliagkas in his role as Divyat Mahajan’s academic advisor, acknowledges support by an NSERC Discovery grant (RGPIN-2019-06512), and a Canada CIFAR AI chair. Pascal Vincent is a CIFAR Associate Fellow in the Learning in Machines & Brains program. Divyat Mahajan acknowledges support via FRQNT doctoral training scholarship for his graduate studies.